OpenStudy (anonymous):

There are n teams in a league and the number of games played is given by the polynomial n^2-n=n. A women's softball league has 14 teams. What is the total number of games to be played?

OpenStudy (anonymous):

28 games to be played

OpenStudy (anonymous):

? what does \[n^2-n=n\] mean

hero (hero):

I was going to say something about that but I wanted someone else to say something first

OpenStudy (anonymous):

166 =14x14

OpenStudy (anonymous):

14 x 14 = 196

OpenStudy (anonymous):

oops 196

OpenStudy (anonymous):

see I'm getting tired now

OpenStudy (anonymous):

maybe it is number of games is \[n^2-n\] in which case since \[n^2-n=n(n-1)\] if n = 14 you get \[14\times 13=182\]

OpenStudy (anonymous):

I came up with 182

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

14P2 ?